OpenStudy (anonymous):

laplace transform of cos (2pi)t

terenzreignz (terenzreignz):

We have to integrate? D:

terenzreignz (terenzreignz):

\[\Large \int\limits_{0}^\infty e^{-st}\cos(2\pi t)dt\]

terenzreignz (terenzreignz):

Come on, @epicChoi make your presence felt :P

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i dont know

terenzreignz (terenzreignz):

You weren't taught how to do Laplace Transforms? D:

OpenStudy (anonymous):

its so fast when it taught to us

terenzreignz (terenzreignz):

Oh... yes, that is a problem... tell you what, let's do it for the general case, where instead of \(2\pi\) we have an arbitrary constant...say, a. And then, when we're done, just replace the a with the constant you desire, in this case \(2\pi\) \[\Large \int\limits_{0}^\infty e^{-st}\cos(2\pi t)dt\]

terenzreignz (terenzreignz):

By the way, THAT is the definition of a Laplace transform... the Laplace transform of a function f(t) is given by \[\Large \int\limits_{0}^\infty e^{-st}f(t)dt\]

terenzreignz (terenzreignz):

So, let's do this.... :) \[\Large \int\limits_{0}^\infty e^{-st}\cos(a t)dt\]

terenzreignz (terenzreignz):

Can you integrate this? Never mind the limits for now, just integrate... \[\Large \int e^{-st}\cos(a t)dt\]

OpenStudy (anonymous):

thanks for lecture

terenzreignz (terenzreignz):

Sarcasm? -.-

OpenStudy (anonymous):

no

OpenStudy (anonymous):

im not good in english :)

terenzreignz (terenzreignz):

Korean, then? I don't speak Korean :P

OpenStudy (anonymous):

hehe

terenzreignz (terenzreignz):

^lol that was random... or not :D

OpenStudy (anonymous):

what is your nationality?

terenzreignz (terenzreignz):

Wouldn't you like to know :D You first :P

OpenStudy (anonymous):

you first please.

terenzreignz (terenzreignz):

We could haggle about who goes first... or we could solve this integral :D \[\Large \int e^{-st}\cos(a t)dt\]

OpenStudy (anonymous):

solve it hehe

terenzreignz (terenzreignz):

But this time, I must insist that you take the first step... at least identify HOW we should integrate this...

OpenStudy (anonymous):

by parts?

OpenStudy (anonymous):

parts or by treating it as the real part of the following complex-valued integral:$$\int e^{(ia-s)t}\,dt$$

terenzreignz (terenzreignz):

Good call :) ... Yikes... this is going to be monstrous... let \(\large u = e^{-st}\) and \(\large dv = \cos(at) dt \) from here, it follows that \(\large du = -se^{-s t}dt\) and \(\Large v = \frac{\sin(at)}{a}\)

OpenStudy (anonymous):

dont be so complicated -_-'

terenzreignz (terenzreignz):

It's going to be complicated one way or another... this is like derivatives... at first, you want to kill yourself (hopefully not really) when you're forced to do it using limits, but eventually, it becomes second-nature due to several rules and useful shortcuts...

terenzreignz (terenzreignz):

So... by parts... \[\Large \int e^{-st}\cos(a t)dt= \frac{e^{-st}\sin(at)}{a}+\int \frac{se^{-st}\sin(at)}{a}dt\]

OpenStudy (anonymous):

please kill me now

terenzreignz (terenzreignz):

Gladly...... LOL what am I saying... oh yeah, do this integral... \[\Large \int e^{-st}\cos(a t)dt= \frac{e^{-st}\sin(at)}{a}+\color{red}{\int \frac{se^{-st}\sin(at)}{a}dt}\]

OpenStudy (anonymous):

i will kill u ?

OpenStudy (anonymous):

thanks for the lecture. i will leave for now. i will buy food for dinner.

OpenStudy (anonymous):

@epicChoi notice \(s/a\) is constant w.r.t. \(t\) hence:$$\int\frac{se^{-st}\sin at}adt=\frac{s}a\int e^{-st}\sin at\,dt$$

terenzreignz (terenzreignz):

Of all the times to do your 'magic' you pick when our OP has left? -.-

OpenStudy (anonymous):

i just want some fried chicken

OpenStudy (anonymous):

give me medal or i tell ur dad on u

OpenStudy (unklerhaukus):

\[\mathcal L\big\{\cos(at+b)\big\}=\frac{s\cos b-a\sin b}{s^2+a^2}\]

OpenStudy (unklerhaukus):

you do not have to integrate, you just have to look it up in a table of laplace transforms, and substitute

OpenStudy (anonymous):

@UnkleRhaukus the whole point is to justify the table's answer :-p

OpenStudy (unklerhaukus):

well that is not clear in the question.

terenzreignz (terenzreignz):

^oh well :)